OpenStudy (anonymous):

Show all work please: Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.

OpenStudy (anonymous):

y-y' = m(x-x') where (x';y')=(-4;2) and m=-1 since the new line is parallel to the old and has same gradient. y-2=-1(x+4) y-2=-x-4 y=-x-2

OpenStudy (anonymous):

the answer is y=-x-2